NON-ADIABATIC PROCESSES IN THE PHOTODISSOCIATION OF QUARTET STATE Na_{3} AND K_{3}

\underline{GIACINTO SCOLES}, JAMES REHO, JOHN P. HIGGINS, CARLO CALLEGARI, AND KEVIN K. LEHMANN, Department of Chemistry, Princeton University, Princeton, NJ 08544.

Excitation of both the 2^{4}E′ ← 1^{4}A′_{2} transition of Na_{3} and the 1^{4}A′′_{2} ← 1^{4}A′_{2} transition of K_{3} lead to curve-crossings into a predissociative doublet state as evidenced by the presence of atomic (P → S) and singlet dimer B → X fluorescing products. Single photon counting techniques have been employed to measure the intersystem crossing times in these simple systems by measuring the onset times of product fluorescence as a function of the vibrational state of the upper quartet state. We find that the spin-orbit mediated intersystem crossing occurs as expected more rapidly for K_{3} than for Na_{3} while, for the latter system, the rapidity of the process increases as we move to higher vibronically excited states.